• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ckeith

    Log entry time 19:13:15 on January15,2002

    Entry number 80339

    keyword=cryotarget control status

    Condition of cryotarget controls as of 1-15-01 at 19:00:
    To the best of my knowledge everything is working ok except:
    1) Fan controls -- low priority since fan settings are rarely adjusted;
    2) Target lifter -- GUI controls are functional; lifter occassionally stops before move is complete. An IOC reboot should fix this. Debug ongoing.
    3) Heater PID -- the CRON job that normally maintains the most up-to-date
    version of the PID settings has been modified to run on the new HallA server. It
    will be started tomorrow.
    4) Alarm limits -- these are no longer automatically written to the IOC bootscript.
    Acceptable values are "hard-coded" for now. I am working with ACE to initiate
    a remote synch between the control computer and server to allow auto update for the
    alarm limits.